### v7

# LLMatic: Neural Architecture Search via Large Language Models and Quality Diversity Optimization

Abstract.
Large Language Models (LLMs) have emerged as powerful tools capable of accomplishing a broad spectrum of tasks. Their abilities span numerous areas, and one area where they have made a significant impact is in the domain of code generation. Here, we propose using the coding abilities of LLMs to introduce meaningful variations to code defining neural networks. Meanwhile, Quality-Diversity (QD) algorithms are known to discover diverse and robust solutions. By merging the code-generating abilities of LLMs with the diversity and robustness of QD solutions, we introduce LLMatic, a Neural Architecture Search (NAS) algorithm. While LLMs struggle to conduct NAS directly through prompts, LLMatic uses a procedural approach, leveraging QD for prompts and network architecture to create diverse and high-performing networks. We test LLMatic on the CIFAR-10 and NAS-bench-201 benchmarks, demonstrating that it can produce competitive networks while evaluating just $2,000$ candidates, even without prior knowledge of the benchmark domain or exposure to any previous top-performing models for the benchmark. The open-sourced code is available in https://github.com/umair-nasir14/LLMatic.
large language models, neural networks, quality-diversity optimization, neural architecture search
journalyear: 2024copyright: acmlicensedconference: Genetic and Evolutionary Computation Conference; July 14–18, 2024; Melbourne, VIC, Australiabooktitle: Genetic and Evolutionary Computation Conference (GECCO ’24), July 14–18, 2024, Melbourne, VIC, Australiadoi: 10.1145/3638529.3654017isbn: 979-8-4007-0494-9/24/07ccs: Computing methodologies Neural networksccs: Computing methodologies Lifelong machine learningccs: Theory of computation Evolutionary algorithms
Figure 1. Illustrated in the figure is the flow of LLMatic. In the initial round of evolution, an initial network with a random prompt goes through a mutation operation. Network individual and prompt individual are then evaluated to be stored in separate archives. During the evolutionary loop, the selected prompt and network go through an evolutionary operation (The prompt is fixed if the operation is crossover) to create more networks and prompt individuals to fill and illuminate the archives.
1. Introduction
A major challenge in deep learning is designing good neural network architectures. Neural Architecture Search (NAS) is the generic term for various approaches to automating this design process (White et al., 2023). The idea is to formulate an objective, such as maximum accuracy on a classification problem with a given budget of parameters and training cycles, and cast the problem as a search for the architecture that maximizes the objective. Every test consists of training the candidate network architecture using some form of gradient descent on the chosen benchmark dataset to measure its performance. This typically means that many thousands of architectures are tested and discarded in the process.
Two common algorithmic approaches to NAS are reinforcement learning and evolutionary computation. Reinforcement learning approaches to NAS (Jaafra et al., 2019) train a controller (typically another neural network) that outputs network architectures; these network architectures are tested and their performance is used as a reward signal. Evolutionary computation approaches to NAS (Liu et al., 2021), on the other hand, directly search the space of neural architectures. A population of architectures are kept, and their performance is used as a fitness score. Evolutionary NAS approaches are similar to neuroevolution, which has existed since the 1980s (Tenorio and Lee, 1988; Miller et al., 1989), and one might even see NAS as a form of neuroevolution. The main difference is that in NAS, the search process does not concern the parameters of the neural network, only its architecture.
One could argue that search by evolutionary computation or reinforcement learning is quite mindless and wasteful, given how many architectures need to be tested and how uninformed the changes that lead to each new architecture are. Is there some way we can inform the search by exploiting stored knowledge about how to design neural networks? This paper explores the idea that we can do exactly this using code-generating large language models (LLMs). More precisely, we propose using LLM with evolutionary algorithm to generate new architectures that have high network architectural diversity and state-of-the-art performance.
The argument for this is simply that modern LLMs fine-tuned on code are very capable (Shinn et al., 2023). Given the amount of machine learning code they have been trained on, it is not surprising that they can design good neural network architectures. However, an LLM by itself cannot, in general, find an optimal architecture for a given problem, as it cannot test architectures and learn from its experiments. Therefore, we propose combining the domain knowledge of code-generating LLMs with a robust search mechanism.
While generating a single architecture that maximizes a given objective is useful for many cases, there is often more value to generating a set of architectures that vary across some relevant dimensions. For example, one might want to have a set of architectures that vary in their parameter counts or depths. This helps in understanding the trade-offs between various desirable metrics and could assist in making better-informed decisions about which architecture to use for a specific application. For example, one might want a range of networks for edge deployments to clients with different RAM sizes. To enable this, the solution we propose here leverages quality-diversity search (Pugh et al., 2016), specifically a version of the MAP-Elites algorithm (Mouret and Clune, 2015).
Our main contribution is a novel LLM-based NAS algorithm, LLMatic, that utilizes the power of two QD archives to search for competitive networks with just $2,000$ evaluations. We empirically show the performance of LLMatic on the CIFAR-10 dataset and the NAS-bench-201 benchmark where LLMatic searches for networks with performance similar to state-of-the-art results.
2. Related Work
Designing neural architectures can be an expensive and unintuitive process for human designers. Neural Architecture Search (NAS) aims to automatically find architectures capable of strong performance after training (Elsken et al., 2019). Bayesian methods are a popular choice given their low sample complexity and the fact that evaluating each architecture (by training it) can be computationally expensive (Kandasamy et al., 2018). Alternatively, reinforcement learning can be used to train an agent (usually another neural network) to output candidate architectures for a given task, with the performance after training of the candidate architecture acting as a reward signal (Jaafra et al., 2019). Evolutionary methods can also be used to search directly through the space of possible architectures (Liu et al., 2021). Similarly, Monte Carlo Tree Search has also been used to search (Wistuba, 2017). In all cases, a human designer must manually define a set of atomic network components or edit actions for use in network search/generation.
To avoid having the designer constrain the space of possible architectures prior to search, we turn to code-generating Large Language Models (LLMs)—large models trained auto-regressively on massive datasets of code (e.g. public repositories hosted on Github). These LLMs are based on the transformer architecture (Vaswani et al., 2017) and have obtained state-of-the-art performance in natural language modelling (Radford et al., 2019; Brown et al., 2020; Adelani et al., 2022; Nasir and Mchechesi, 2022). They have also been successfully used in specific applications, such as for video game level design (Todd et al., 2023; Nasir and Togelius, 2023; Chen et al., 2021) or code generation (Shinn et al., 2023).
Recently, LLMs have been used for evolving code by framing code-generation as an evolutionary problem. Evolution through Large Models (ELM) (Lehman et al., 2022) casts LLMs as evolutionary operators within a MAP-Elites (Mouret and Clune, 2015) algorithm tasked with evolving a robot’s morphology at code-level. EvoPrompting (Chen et al., 2023) is an LLM-based method that is somewhat similar to ours in that it uses code-LLMs as mutation and crossover operators to perform NAS. It is tested on the MNIST-1D classification task (Greydanus, 2020) and the CLRS algorithmic reasoning benchmark (Veličković et al., 2022). Since performance can generally be trivially increased by simply adding parameters to the model, an additional penalty is added to the fitness of a candidate neural architecture corresponding to its model size. This incentivizes the discovery of small models with effective architectures. In our method, we instead consider model complexity (in terms of FLOPS) as a diversity metric, searching for high-performing models of a variety of sizes. GENIUS (Zheng et al., 2023) is another LLM-based NAS algorithm that uses GPT-4 to simply search through straight-forward prompting.
Quality Diversity (QD) methods (Pugh et al., 2016) are a family of evolutionary algorithms that, in addition to optimizing a fitness metric, search for a diversity of individuals according to some user-specified “behavioral descriptors”. Instead of keeping a population of the fittest individuals, QD methods such as MAP-Elites (Mouret and Clune, 2015) maintain an “archive” of individuals, where this archive is partitioned into cells, with each cell corresponding to individuals exhibiting a particular range of values along each behavioral descriptor.
QD methods are valuable in domains such as robot control, where it is useful to learn diverse high-quality trajectories, in case one solution should become unavailable during deployment because of a physical obstruction or mechanical malfunction (Cully et al., 2015). Another motivating factor is that greedily searching for the fittest individual may not be desirable in deceptive domains. Here, maintaining a diversity of fit individuals may protect the population from falling into local optima (Gaier et al., 2019). Conversely, diverse, unorthodox solutions may provide valuable “stepping stones” on the path to globally fit individuals.
3. Approach
1  Initialize network and prompt archives. while Number of generations $<$ maximum generations
2       foreach Network in batch of networks do
3             if Number of individuals in archives $<$ set threshold then
4                   Mutate the initial network with a random prompt. Get the network and prompt individuals and add them to the respective archives.
5            else
6                   randomly choose mutation or crossover as the evolutionary operator. if evolutionary operator ==absent$=$ = mutation then
7                         Select network and prompt. Mutate the selected network with the selected prompt. Train or query the generated network. Get the network and prompt individuals and store them.
8                  else
9                         Selection of networks for crossover. Perform crossover on selected networks with a fixed prompt. Train or query the generated network. Store the generated network individual.
10                   end if
11
12             end if
13
14       end foreach
15      Evaluate all networks to find the test accuracies. Mutate the temperature for the next generation. Add the batch network individuals and the corresponding prompt individuals in the respective archives.
16 end
Algorithm 1 LLMatic
LLMatic begins its search with a very basic neural network, inspired by the work of  (Stanley and Miikkulainen, 2002) which suggests that neuroevolution tends to perform better when starting with a small network. In LLMatic, we use a novel dual-archive cooperative QD optimization approach, in which two separate archives are used to store complementary components that can be combined to solve a given task. The first archive stores neural networks, where the width-to-depth ratio and Floating Point Operations per Second (FLOPS) of a network are the behavioural descriptors. The width-to-depth ratio is a division of the width and the depth of the network. To specify the width, we use the maximum of the output features of all layers, while depth is simply the number of layers. Note that we choose FLOPS instead of parameter count because FLOPS correlates better with actual time spent training a network (Ayala et al., 2017). We call this archive the “network archive”. The fitness function for the networks in this archive is defined as the test accuracy of the network after training. The second archive, called the “prompt archive”, contains the prompt and temperature used for generating code, which are also the behavioural descriptors. Temperature of an LLM is a parameter that controls the creativity of the LLM. Lower temperature means the LLM will not diverge from the answer, higher means it will be more creative. The selection of prompt and temperature depends on a curiosity score (Cully and Demiris, 2017), governed by whether the generated network was added to the network archive. The fitness of prompt individuals depends on whether the network was better than the previous generation’s best score. Figure 1 illustrates the flowchart of the approach LLMatic uses.
In the first generation, a simple neural network with one convolutional and one fully connected layer initiates the evolution [line 1 of algorithm 1]. A prompt is selected at random to generate an initial batch of networks [line 5-6 of algorithm 1]. These networks are evaluated and an attempt is made to add them to the network archive as a random initialization for MAP-Elites. Concurrently, we mutate the temperature based on the fitness of the network, increasing it if the fitness increases and vice versa [line 21-23 of algorithm 1]. By increasing the temperature, we want the LLM to explore as it is performing better than before. By decreasing the temperature, we want the LLM to exploit and try to achieve better fitness than before. Once we calculate the fitness of the prompt individual, we add the score to a collective prompt fitness score, after which we try to populate the prompt archive. The collective prompt fitness score determines the overall fitness of each individual in the prompt archive as it gives each prompt a fitness score.
Once either of the archives reaches a specified capacity, we introduce neural network training and evolutionary operators in the process [line 7-20 of algorithm 1]. With a certain probability at each generation, a decision is made on whether to perform crossover or mutation to produce $N$ new offspring. If the crossover operator is chosen, we select $N$ random network individuals, locate their closest networks in the archive, and carry out a crossover operation instructed by a prompt [line 15-16 of algorithm 1]. No individual is added to the prompt archive when a crossover is performed. If the mutation operation is selected, we pick the most curious prompt individual and a random network individual. For exploration, we also select random prompts. In both cases, each network is trained for a certain number of epochs and an attempt is made to add the network to the archive. Likewise, a prompt individual is added as previously described. This process continues for a pre-determined number of generations. Algorithm 1 shows the complete search process of LLMatic in pseudocode. Refer to supplementary material for pseudocode on mutation operators, crossover operators, temperature mutation and addition to archives.
4. Evaluating LLMatic
To evaluate LLMatic, we use CIFAR-10 (Krizhevsky et al., 2009), a commonly used dataset for NAS (Tan and Le, 2019; Ying et al., 2019). We perform extensive ablation studies to demonstrate that LLMatic benefits from each of its components during search. Once our algorithm is validated, we extend our experiments to NAS-bench-201. The NAS-bench-201 benchmark (Mehta et al., 2022) is a dataset enumerating all possible neural architectures within a given search space (i.e. of a fixed number of nodes/layers and edges/operations) and the corresponding test accuracy of each architecture after training on a given dataset (e.g. CIFAR-10), allowing researchers to explore the tradeoffs of various NAS algorithms without needing to retrain each candidate network during search.
4.1. Setting up LLMatic
Dataset: The CIFAR-10 dataset is made up of $60,000$ color images, each with a resolution of $32\times 32$ pixels, and divided into 10 categories: airplane, automobile, bird, cat, deer, dog, frog, horse, ship and truck. The dataset is partitioned into five groups for training and one group for testing, each group holding 10,000 images. Each test group consists of an exact count of 1,000 images from each category selected randomly. The training groups hold the remaining images, which are arranged in random order. As a result, some training groups might contain more images from one category compared to others. Nonetheless, collectively, the training groups have an exact total of 5,000 images from each category.
Initial Neural Network: LLMatic starts off with a simple neural network with one convolutional layer that takes in $3$ input with 3 channels, with $1\times 1$ kernel size and $1$ output channel which connects to a dense layer of size $1024$. These hidden neurons are connected via another dense layer to $10$ output neurons (as we have 10 classes). Rectified Linear Unit (ReLU) (Nair and Hinton, 2010) is the activation function used in all layers. All of our networks are generated in PyTorch (Paszke et al., 2019).
Generating Neural Networks: At each generation, we generate a batch of $100$ new offspring. Each network generated is trained for $50$ epochs. The networks are optimized by stochastic gradient descent (Bottou, 2010) with the learning rate set to $0.001$ and momentum set at $0.9$ for all networks. We use cross entropy loss as our measure for the fitness of the trained network.
Figure 2. The illustration of the best accuracy per generation for LLMatic and all ablation studies. Each experiment is conducted with 30 seeds. The shaded region is the standard deviation while the solid line represents the mean. EfficientNet-B0 is the best-performing EfficientNet on CIFAR-10.
For evolutionary operators, we set a probability of $0.7$ for mutation and $0.3$ for crossover as after experimentation, we found that mutation creates consistently more trainable neural networks. We initialize the temperature parameter (used when sampling the code-generating LLM) to $0.6$. For temperature mutation, half of the population is generated by the prompt individual temperature mutated uniformly at random between $-0.1$ to $0.1$. The other half is generated by the temperature obtained from the prompt individual itself. If the fitness of the generated network is better than or equal to the best fitness of the previous generation, we increase the temperature by $0.05$ and if it is worse than the best fitness of the previous generation, we decrease it by $0.05$. For the crossover operator, we select $10$ random networks and find their $2$ or $3$ nearest neighbours in the network archive to perform crossover. We set the temperature to be $0.7$ for network generation.
Quality Diversity Optimization: For our QD optimization algorithm, we choose a variant of MAP-Elites, called Centroidal Voronoi Tessellation (CVT-MAP-Elites) (Vassiliades et al., 2017) can be seen as a generalization of MAP-Elites, intended to ease scaling up MAP-Elites to high-dimensional behavior spaces; CVT-MAP-Elites outperforms standard MAP-Elites in high-dimensional spaces while matching performance lower-dimensional scenarios such as those studied in the present work. CVT-MAP-Elites automates the sub-division of the archive by identifying $k$ cell centroid locations exhibiting an even spread through the behavioural descriptors space. (Here, $k$ corresponds to the total number of evolutionary “niches” in the archive.) We use the pymap_elites [^1] https://github.com/resibots/pymap_elites implementation for our experimentation. k-d tree (Bentley, 1975) is used to create and write centroids to the archive and find the nearest neighbors using a Euclidean distance metric (Dokmanic et al., 2015).
Each of our QD archives has $2$ dimensions (behavioral descriptors), with $100$ niches spread across them. We set our number of random initial networks to $10$. For the network archive, we have the width-to-depth ratio of the network as our first dimension and the FLOPS of the network as the second dimension. The width-to-depth ratio has a lower limit of 00 and an upper limit of $200$. The minimum FLOPS is set to $200$ MegaFLOPS and the maximum is set to $5$ GigaFLOPS. This range is set after experimentation.
For the prompt archive, we have the prompt encoded as an integer as the first dimension and temperature as the second dimension. The maximum value of the prompt is 16, the number of prompts used in the system. The maximum temperature value is set to $1$ as it can never increase beyond that for our LLM. The lower limit for all dimensions is 00.
For the network archive, we simply select a random network while for the prompt archive, we select the most curious prompt individual, which depends on the curiosity score. This curiosity score is incremented by $1.0$ if the selected prompt adds the generated network to the network archive, decreased by $0.5$ if the network is not added, and reduced by $1.0$ if the created network is untrainable. If the generated network has better fitness than the previous generation’s best network, the collective prompt fitness score for the prompt in the prompt individual is increased by $1$, otherwise, it is unchanged. We use prompts that are generalizable to any problem in any domain. Refer to LABEL:appendixprompts for an example of mutation and crossover prompts.
Figure 3. The illustration of how many trainable networks are created in a generation. The total number of networks created is 100 per generation.
Code Generating LLM: We use the pre-trained CodeGen (Nijkamp et al., 2022) LLM to generate neural networks. CodeGen is an autoregressive decoder-only transformer with left-to-right causal masking. CodeGen is first trained on ThePile dataset with random initialization and is called CodeGen-NL. CodeGen-Multi is initialized with CodeGen-NL and is trained on BigQuery dataset. Lastly, CodeGen-Mono is initialized with CodeGen-Multi and is trained on BigPython. CodeGen is trained to be in various parameter sizes but we use $6.1$ Billion parameter variant of CodeGen-Mono due to computational constraints.
ThePile dataset (Gao et al., 2020) is an $825.18$ GB English text corpus. (Nijkamp et al., 2022) select a subset of the Google BigQuery dataset which contains 6 programming languages, namely C, C++, Go, Java, JavaScript, and Python. The authors collected a large amount of permissively licensed Python code from GitHub in October $2021$, and named it BigPython. The size of BigPython is $217.3$ GB.
CodeGen-6B has $33$ layers and $16$ heads with $256$ dimensions per head. The context length is $2048$ and the batch size is $2$ million tokens. Weight decay is set to $0.1$. $0.4e^{-4}$ is the learning rate. Warm-up steps are set to $3k$ while total steps for training are $150$k.
4.2. Ablation Study
As we have many components in LLMatic, we choose to do a thorough ablation study to determine the effect of each component on overall performance. The following are the components tested for the ablation study:
  • Network-Archive-LLMatic: LLMatic with only the network archive. To achieve this, we create a population of prompt individuals. The population is fixed to $100$ individuals initialized with random individuals. We have only one fitness score for this population, which is calculated as $+1$ if a network is added in the network archive, $-0.5$ if the network is not added and $-1$ if the network is not trainable. After we generate the network, we mutate temperature by adding $0.1$ if the network is added in the network archive and $-0.1$ if the network is not added.
  • Prompt-Archive-LLMatic: LLMatic with only the prompt archive. To achieve this, we create a population of networks. The fitness function for the population of networks is accuracy. We keep the population to $100$ individuals. With a similar probability as LLMatic, we select mutation or crossover operator. For the crossover operator, we select the individual that is closest to the structure of the selected network. For network similarity, we use cosine similarity and we choose the networks with higher scores. For the mutation operator, similar to LLMatic we mutate half of the networks from the most curious prompt individuals and half from random individuals.
  • Mutation-Only-LLMatic: LLMatic using only mutation.
  • Crossover-Only-LLMatic: LLMatic using only crossover.
  • Random-NN-Generation: Neural network generation without evolution. We generate $100$ networks per generation for $20$ generations as a fair comparison to LLMatic, which generates the same number per batch. We apply the prompt “Create a neural network that inherits from nn.Module and performs better than the above neural network” and we add the initial network with this prompt.
4.2.1. Ablation Results and Discussion
In this section, we will discuss the results of the experiments that we set up in the previous section. We first discuss the best loss per generation, illustrated in Figure 2. This will lead our discussion to trainable networks generated by changing the crossover and mutation probabilities (Figure 3). Then we will discuss how archives are illuminated Figure 4. Some of the generated networks are shown in the supplementary material.
(a) Prompt archive: Prompts encoded as integers on the $x$-axis, normalised to be in range 0-1 for CVT-MAP-Elites as all points are within 001111. On the $y$-axis, we have the temperature that controls LLMs exploration ability. As 1 is the maximum temperature, there is no need for normalisation.
(b) Network archive: Width-to-depth ratio on the $x$-axis. The range for the Width-To-Depth ratio is from 00200200200200 normalised to 001111. On the $y$-axis, we have Floating Point Operations per Second (FLOPS). We have a range of 200 Mega FLOPS to 5Giga FLOPS. This range is normalised to 0-1 for CVT-MAP-Elites.
Figure 4. An illustration of archives generated by LLMatic. We have selected the archive with the median number of cells filled in experiments over 30 seeds. 3(a) shows prompt archive. 3(b) shows network archive. The lighter the colour of the filled cell, the better fitness of the individual.
Looking at Figure 2, it is clear that each component of LLMatic is necessary. Mutation-Only-LLMatic and Network-Archive-LLMatic are the closest to LLMatic which also indicates that our choice of giving mutation more probability of being selected is the right one. Crossover-Only-LLMatic is understandably the worse as mutation provides more exploration (Ullah et al., 2022). Both operators, mutation and crossover, together give exploration and exploitation abilities to LLMatic, which are highly necessary to find high-quality and diverse networks. While Prompt-Archive-LLMatic does significantly worse as network archive is an important aspect to find high-performing networks. Both archives together demonstrate competitive results.
We use EfficientNet-B0, which is the state-of-the-art network on CIFAR-10 Tan and Le (2019) as an indicator of where our algorithm stands. EfficientNet-B0 was searched via methods applied by Tan et al. (2019) and is slightly larger than the original study as they were targeting more FLOPS. The original study required $8,000$, while LLMatic requires $2,000$ evaluations to find a competitive network. EfficientNet-B0 was first trained on the ImageNet dataset (Deng et al., 2009) and then on CIFAR-10 via transfer learning (Torrey and Shavlik, 2010). This is an advantage for EfficientNet-B0 as ImageNet has many classes and is an order of magnitude larger dataset.
Figure 4 demonstrates how each archive is filled on average. We can see the prompt archive contains high-performing individuals who have the first few prompts and higher temperatures.
Some of the high-performing individuals do have lower temperatures, which suggests that sometimes it is good to pick deterministic layers. For network archives, we observe a diversity of high-performing networks with respect to both FLOPS and width-to-depth ratio. More than $20$ networks are competitive networks in this archive.
To delve into why we choose the probabilities being $0.7$ for mutation and $0.3$ for crossover, we observe the number of trainable networks generated per generation (see Figure 3). This is to be considered common knowledge that the more working individuals we have, the greater the chance of high-performing individuals. For this purpose, we train LLMatic with uniform probabilities, and $0.3$ for mutation and $0.7$ for crossover. We observe that uniform probabilities are still competitive with the original setting, while increasing the crossover probability makes it worse. The results of these experiments and results of the ablation study for Mutation-Only-LLMatic and Crossover-Only-LLMatic lead us to the conclusion that mutation should be given more probability of being selected.
5. Experiments on NAS-bench-201
(a) CIFAR-10
(b) CIFAR-100
(c) ImageNet16-120
Figure 5. Illustration of test accuracies of all networks across all datasets and best-found networks in each generation by LLMatic.
Next, we extend our experimentation of LLMatic on NAS-bench-201 benchmark (Dong and Yang, 2020), which searches a cell block for a constant neural network structure. The structure is initiated with one 3-by-3 convolution with 16 output channels and a batch normalization layer (Ioffe and Szegedy, 2015). The main body of the skeleton includes three stacks of cells, connected by a residual block. Each cell is stacked 5 times, with the number of output channels as 16, 32 and 64 for the first, second and third stages, respectively. The intermediate residual block is the basic residual block with a stride of 2 (He et al., 2016), which serves to downsample the spatial size and double the channels of an input feature map. The shortcut path in this residual block consists of a 2-by-2 average pooling layer with a stride of 2 and a 1-by-1 convolution. The skeleton ends with a global average pooling layer to flatten the feature map into a feature vector. Classification uses a fully connected layer with a softmax layer to transform the feature vector into the final prediction.
The specified cell within the search domain is depicted as a densely connected directed acyclic graph with four nodes and six edges; here, nodes symbolise feature maps while edges denote operations. There are five possible operations: (1) zeroize, (2) skip connection, (3) 1-by-1 convolution, (4) 3-by-3 convolution, and (5) 3-by-3 average pooling layer. Zeroize drops out the associated edge operation. Given five operations to choose from, the aggregate count of potential search spaces comes to $156255^{6}=$ 15625 cell combinations. Evaluations are carried out on CIFAR10, CIFAR100 (Krizhevsky et al., 2009), and ImageNet16-120 (Chrabaszcz et al., 2017). ImageNet16-120 is a variant of ImageNet dataset (Russakovsky et al., 2015) which is downsampled to 16x16 image sizes and contains the first 120 classes.
5.1. Results
To remain consistent with our previous experiments, LLMatic searches for $20$ generations and $100$ cells in a generation. We curate the prompt to cater for a controllable generation by restricting it to the five operations. Refer to supplementary material for an example of how we generate queryable cells. For our network archive, we take minimum and maximum FLOPS as the bounds for the behaviour descriptor.
Table 1. A comparison of test accuracy on NAS-bench-201 LLMatic with $\Lambda$-DARTS (Movahedi et al., 2022), having the most near-to-optimal result, GENIUS (Zheng et al., 2023), a GPT-4 based neural architecture search algorithm, Random Search implemented by NAS-bench-201 authors (Dong and Yang, 2020) and DARTS (Liu et al., 2018). We provide optimal accuracy for reference, which is the maximum accuracy that can be achieved in NAS-bench-201. LLMatic achieves near-to-optimal results but has more variation than the comparing methods.
Method CIFAR-10 CIFAR-100 ImageNet16-120
DARTS 54.30±0.00 15.61±0.00 16.32±0.00
Random Search 93.70±0.36 71.04±1.07 44.57±1.25
GENIUS 93.79±0.09 70.91±0.72 44.96±1.02
$\Lambda$-DARTS 94.36±0.00 73.51±0.00 46.34±0.00
LLMatic 94.26±0.13 71.62±1.73 45.87±0.96
Optimal 94.47 74.17 47.33
We compare our results with GPT-4 based NAS algorithm GENIUS (Zheng et al., 2023) as an LLM baseline, $\Lambda$-DARTS (Movahedi et al., 2022) as it achieves close to optimal result, where optimal is the maximum test accuracy, and Random Search. As Table 1 indicates, LLMatic has better results than simple GPT-4 based NAS and close to the state-of-the-art and optimal results.
Furthermore, in Figure 5 we look into the found networks by LLMatic over each generation. We observe that the distribution of found networks is spread wide in the search space. This is due to the procedural nature and exploration capabilities of LLMatic through the prompt archive. To demonstrate near-to-optimal networks we look into Table 2 for maximum ranked networks based on test accuracies searched by LLMatic.
Table 2. Maximum rank achieved by LLMatic on each dataset in NAS-bench-201.
Method Rank
CIFAR-10 2
CIFAR-100 2
ImageNet16-120 11
6. Conclusion and Future Work
To conclude, we present LLMatic: a novel Neural Architecture Search (NAS) algorithm that harnesses the power of Large Language Models (LLMs) and Quality-Diversity (QD) optimization algorithms. LLMatic successfully finds competitive networks that are diverse in architecture. We show empirically that LLMatic can find more than 20 competitive networks in CIFAR-10 and near-to-optimal networks in NAS-bench-201, using only 2000 evaluations. LLMatic decreases the max population size per generation to only 100. LLMatic achieves this while relying on a 6.1B parameter language model. Furthermore, we show that each component in LLMatic is necessary. We do an extensive ablation study and find that LLMatic finds the network with the best accuracy among other variants.
LLMatic achieves this with many constraints in hand. Firstly, we use CodeGen-6.1B code generation LLM, which is a smaller language model when compared to existing LLMs. This demonstrates how computationally efficient LLMatic is, and gives us reason to believe that further gains could be unlocked by incorporating larger language models Secondly, due to computational resources, we keep our searches to $2000$, and still find competitive networks.
In future work, LLMatic should be compared to other NAS methods on other computer vision and natural language processing tasks. As neuroevolution is similar to NAS, LLMatic could be compared to reinforcement learning benchmarks as well. With this, LLMatic can be used in tasks such as open-ended learning as well (Nasir et al., 2022).